Solve the given system of differential equations by systematic elimination.
Dx + D2y = e3t
(D + 1)x + (D 1)y = 4e3t
Answer #1

Solution:

Dr +Dy

3t

Multiply equation \small \left ( i \right ) by \small \left ( D+1 \right )    and equation \small \left ( ii \right )    by \small D    and subtract the resulting equations

\small D\left ( D+1 \right )x+D^{2}\left ( D+1 \right )y=\left ( D+1 \right )e^{3t}

\small D\left ( D+1 \right )x+D\left ( D-1 \right )y=4De^{3t}

\small -                  \small -                         \small -

------------------------------------------------------------

\small 0x+\left ( D^{3}+D \right )y=\left ( D+1 \right )e^{3t}-4De^{3t}

\small \therefore \left ( D^{3}+D \right )y=3e^{3t}+e^{3t}-12e^{3t}=-8e^{3t}

The Auxiliary equation is \small D^{3}+D=0

\small D\left ( D^{2} +1\right )=0\Rightarrow D=0,\pm i

The complementary function is   \small y_{c}=c_{1}+c_{2}cost+c_{3}sint

Particular integral \small \left ( y_{p} \right )

\small =\frac{1}{f\left ( D \right )}\left ( -8e^{3t} \right )

\small =\frac{-8}{D^{3}+D}\left ( e^{3t} \right )

\small f\left ( D \right )=D^{3}+D

\small \therefore f\left ( 3 \right )=3^{3}+3=27+3=30\neq 0

\small \therefore y_{p}=\frac{-8}{30}e^{3t}=\frac{-4}{15}e^{3t}

\small \therefore y=y_{c}+y_{p}

\small \therefore y=c_{1}+c_{2}cost+c_{3}sint-\frac{4}{15}e^{3t}

Substitute    \small y=c_{1}+c_{2}cost+c_{3}sint-\frac{4}{15}e^{3t} in equation \small \left ( i \right )

\small Dx+D^{2}\left ( c_{1}+c_{2}cost+c_{3}sint-\frac{4}{15}e^{3t}\right )=e^{3t}

\small \therefore Dx-c_{2}cost-c_{3}sint-\frac{12}{5}e^{3t}=e^{3t}

\small \therefore Dx=c_{2}cost+c_{3}sint+\frac{17}{5}e^{3t}

\small \therefore x=\int \left ( c_{2}cost+c_{3}sint+\frac{17}{5}e^{3t} \right )dt

\small \therefore x= c_{2}sint-c_{3}cost+\frac{17}{15}e^{3t}

\small \thereforeThe solution to the system are

\small x= c_{2}sint-c_{3}cost+\frac{17}{15}e^{3t}\: ,\: \: \: \small y=c_{1}+c_{2}cost+c_{3}sint-\frac{4}{15}e^{3t}